OpenStudy (henryarias14):

CALCULUS - Derivatve of natural log i need to find the derivzative of ln x(3x-2). i know its the chain rule but i cant figure it out; this is what i have so far:

zepdrix (zepdrix):

Is everything inside of that log?\[\large\rm \ln\left[x(3x-2)\right]\]Like that?

OpenStudy (henryarias14):

i got the outside function which is \[\frac{ 1 }{ x(3x-2) }\]

OpenStudy (henryarias14):

yes

OpenStudy (henryarias14):

but i dont know if i have to treat x as g(x) and (3x-2) as h(x) in the chain rule. idk i got stuck there, this is part of a bigger natural log

zepdrix (zepdrix):

Ah ok good. First step looks good. We'll chain rule into product rule. It might feel a little weird at first.\[\large\rm \frac{d}{dx}\ln\left[x(3x-2)\right]\quad=\quad \frac{1}{\left[x(3x-2)\right]}\left[x(3x-2)\right]'\]

OpenStudy (henryarias14):

oh i see, alright

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\ln\left[x(3x-2)\right]\quad=\quad \frac{1}{\left[x(3x-2)\right]}\left[x'(3x-2)+x(3x-2)'\right]\]

OpenStudy (tehhota):

I would simplify the (x(3x-2)) So you dont have to use the product rule

OpenStudy (tehhota):

Saves you an extra step

OpenStudy (henryarias14):

oh thats an awsome idea tehHota. hadnt think of that

zepdrix (zepdrix):

Yes, simplifying the inside of the log before differentiation is a good option. Or expanding the expression into two separate log using log rule is another option as well.

OpenStudy (henryarias14):

how would you separate it? x and (3x-2) right?

OpenStudy (tehhota):

Sometimes you cant simplify the inside so its always good practice to do it the long way

OpenStudy (henryarias14):

yeah i know

OpenStudy (tehhota):

Yes

OpenStudy (tehhota):

x and (3x-2)

OpenStudy (henryarias14):

ok, lets do it without simplifying for practice. got a final in a week

OpenStudy (henryarias14):

with the product rule

zepdrix (zepdrix):

Oh boy :D fun stuff

OpenStudy (tehhota):

I prefer anti-derivatives

OpenStudy (henryarias14):

yeah lol fun stuff and this is only part of a BIGGER problem lol. trying to logarithmic differentiation on \[\frac{ x^2 \sqrt{3x-2}}{ (x+1)^2 }\]

OpenStudy (henryarias14):

thats a nasdty problem lol

OpenStudy (henryarias14):

ok so for the product rule i got this

OpenStudy (henryarias14):

\[\frac{ 1 }{ x(3x-2) } * (6x-2)\]

OpenStudy (henryarias14):

please check it, im not sure if its right or not

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\ln\left[x(3x-2)\right]\quad=\quad \frac{1}{\left[x(3x-2)\right]}\left[x'(3x-2)+x(3x-2)'\right]\]\[\large\rm \frac{d}{dx}\ln\left[x(3x-2)\right]\quad=\quad \frac{1}{\left[x(3x-2)\right]}\left[1(3x-2)+x(3)\right]\]\[\large\rm \frac{d}{dx}\ln\left[x(3x-2)\right]\quad=\quad \frac{1}{\left[x(3x-2)\right]}\left[6x-2\right]\]Mmm ya looks correct.

OpenStudy (tehhota):

Confirmed

OpenStudy (henryarias14):

thank you!! this is my last math before i transfer schools. not a big fan of calculus

OpenStudy (tehhota):

It can get confusing

OpenStudy (tehhota):

I hated the chain rule

zepdrix (zepdrix):

WHUT Calculus is the best!! :D Ya chain rule is definitely the hardest of the simple differentiation rules. Takes some practice.

OpenStudy (henryarias14):

its just hard to see the different functions inside the chain rule. its hard when you have to combine it with product rule and others

OpenStudy (henryarias14):

one last question. i could simplify the last answer like this right? \[\frac{ 6x-2 }{ x(3x-2) }\]

OpenStudy (tehhota):

Technically even with simple \[5x^2\] equations you still use the chain rule, but it only multiplies by 1 so we leave that part out

zepdrix (zepdrix):

Yes. You could even factor a 2 out of the terms in the numerator, but that's unnecessary. The way you wrote it is preferable.

OpenStudy (tehhota):

The only good part of calculus, leaving things expanded.

OpenStudy (henryarias14):

alright, thanks

OpenStudy (henryarias14):

yeah my professor dont mind if i dont simply to the end

OpenStudy (xapproachesinfinity):

Another way is to make use of log properties first before differentiating i.e \[\ln [x(3x-2)]=\ln x+\ln (3x-2)\] now we launch into derivative \[\frac{d}{dx}[\ln(x(3x-2))]=\frac{d}{dx}[\ln x]+\frac{d}{dx}[\ln (3x-2) ]\]

OpenStudy (xapproachesinfinity):

just an alternative if you like